OpenStudy (anonymous):

limit of x---> infinity (1/e^x)????please help

OpenStudy (anonymous):

0

OpenStudy (anonymous):

please explain

OpenStudy (anonymous):

As x goes to infinity e^x goes to infinity right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So what is 1/infinity? or, 1/(A REALLY LARGE NUMBER) It approaches 0.

OpenStudy (anonymous):

i saw that min x--->-infinity (e^x)=0 is it something similar

OpenStudy (anonymous):

even though for this its -infinity?

OpenStudy (anonymous):

1/10 = .1 1/100 = .01 1/1000 = .001 1/10000 = .0001 1/100000 = .00001 Let the denominator go to 10^10^10^10 or something haha Well if you have: \[\lim_{x \rightarrow \infty} \frac{1}{e^x} = \lim_{x \rightarrow \infty} e^{-x}\] Just let x -> -x \[\lim_{x \rightarrow \infty} e^{-x} =\lim_{-x \rightarrow \infty} e^{x}=\lim_{x \rightarrow -\infty}e^{x} = 0\]

OpenStudy (anonymous):

ohh wow i see okay so it works for both pos or neg infinity as im understanding

OpenStudy (anonymous):

Well, since one is 1/e^x and the other is e^x.

OpenStudy (anonymous):

If you had: \[\lim_{x \rightarrow \infty} e^x \rightarrow \infty\]

OpenStudy (anonymous):

but it was neg infinity it would be 0?

OpenStudy (anonymous):

*if

OpenStudy (anonymous):

Thank you!